A Complete Bibliography of Publications in Synthese, 1940–1949

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.02

Title word cross-reference

16th [134, 106]. 1848 [80]. 1907 [107]. 1941 [54]. 1946 [106]. 1st [54].

Académie [93]. Academy [119, 48, 131, 93]. accuracy [165]. Actes [140]. activity [26, 231]. Address [134, 38, 59, 106, 211, 192]. After [27]. Afternoon [54]. against [218, 117, 208]. Algemeen [37]. Ambassade [39]. Ambassador [14]. Amsterdam [33, 134]. analysis [204]. ancien [179]. année [76]. Annual [74]. April [113]. architecture [219]. aspect [212]. Aspecten [99, 52]. Aspects [53, 111, 152, 163, 99, 52]. Aspekte [163]. Assemblée [224]. Assembly [224]. Aufforderungsfunktion [86]. August [125]. aux [39]. Avant [77]. Avant-propos [77]. avril [113].

B [199]. B. [103]. Bas [39]. basic [24]. basis [186]. be [93]. Bedeutung [128]. Bergson [70]. Berne [113, 113, 100]. better [25]. between [69]. Bibliography [69]. Bij [74]. biological [202]. Biologie [52, 160, 116]. Biology [52, 53, 160, 116]. biophysics [203]. Boekbespreking [11].

Boerma [228]. Book [21, 124, 136, 143, 135, 217, 191, 226, 168, 103, 32, 96, 11]. Book-Committee [124, 136]. books [51, 120, 169]. Boole [144]. Bridgman [218]. British [56]. Brouwer [134, 107, 106]. Burgomaster [193]. buts [90].

C [199, 199]. Cahiers [43]. calculi [31]. calculus [198]. Carnap [229]. ce [214]. centenary [144]. Center [121, 42, 171]. Centre [171, 42, 19, 121]. champ [26]. changes [24]. Channels [154]. chapters [189]. character [123]. Ching [189]. Christian [123]. Clay [182, 183]. co [13, 153]. co-operation [13, 153]. Collective [206]. Combination [230]. Commemoration [107]. Committee [124, 136]. Communication [121, 121]. comprehensibility [201]. comprehension [25]. concept [181]. concepts [204]. concernant [82]. concerning [82, 137]. Concluding [61]. conditions [133, 180]. Conference [57, 125, 156, 157, 157, 155]. conferment [134]. Congres [233, 140, 113, 100]. Congress [233, 140, 113, 100]. consideration [111]. Constructions [62]. contact [73]. Contest [220]. Contingency [190]. contre [117]. control [132]. Cooperation [154, 154]. cornerstones [227]. cosmic [183]. cours [93]. course [7]. crises [49]. current [93]. cybernetics [203]. cycles [49].

d’activité [26]. dans [78, 154, 101, 89, 117, 114]. day [107]. defence [209]. definitions [72, 87]. degree [134]. delivered [134, 106]. depth [128]. d’été [157]. d’études [42, 171]. development [68, 137]. developments [213]. Dienste [91]. Different [196]. Différents [196]. d’intersubjectivité [196]. d’introduction [6]. Diplomacy [151]. Diplomatie [151]. discourse [79]. Discussion [31]. Doctor [134]. doctorate [106]. dogma [123]. Dr [107]. droit [80]. during [41, 93]. Dutch [74, 26, 1, 34, 99, 112, 9, 11].

economic [68]. Economics [15, 19]. economists [49]. Economy [188]. Editorial [105]. Editors [33]. eight [18]. Einheit [160]. Elections [131]. Embassy [39]. Empire [9]. empirical [207, 98]. Empiricism [172, 173, 227, 186]. Empirie [9]. empirique [207]. English [134, 35, 20]. Ernest [31]. Erratum [88]. errors [29]. Esprit [62, 114]. Essay [220, 175]. Essentials [110]. ethics [9, 162, 209]. ethiek [9]. étude [194, 26]. étudiés [93]. Everyday [66]. evidence [82, 82]. evolution [181]. Exaktheit [165]. Excellency [14]. Exchange [74]. experience [215].

fact [161]. Falsification [29]. February [54, 107]. field [26]. fight [117, 208]. Final [211]. Fixed [132]. fondation [121]. Foreword [77]. Formal [146, 231, 147]. formalisée [147]. formalisierter [201]. formalism [218]. formalized [201]. Forschung [128]. foundation [108, 229, 121]. foundations [216, 28]. Fourth [156, 157]. Francaise [40, 232, 41, 140]. France [39, 80]. freedom [164]. French [140, 77, 26, 43, 6, 76, 25, 40, 78, 113, 100, 121, 171, 157, 154, 233, 82, 140, 80, 151, 207, 101, 195, 148, 200, 232, 214, 147, 90, 39, 180, 42, 179, 224, 205, 194, 62, 127, 41, 89, 117, 196, 187, 197, 122, 93, 92, 114]. function [86].

G [134]. G. [106]. Gedanken [199]. gedeelte [37]. Geleide [112, 2]. General [64, 98, 37, 192]. générale [37]. genetics [86, 198]. German [201, 165, 199, 160, 115, 52, 116, 86, 91, 222, 30, 198, 163, 128, 178, 188, 223]. geschichtliche [128]. Gesellschaft [178, 223]. Ginneken [10]. Goals [90]. Gradation [202]. graduation [107]. Graveland [125, 193]. greeting [193]. Group [26]. Groupe [26]. growth [145]. guerre [180, 41]. Guided [112]. guilt [206].

Hague [14]. happiness [133]. hedonism [225]. held [54]. Henri [70]. Heymans [110]. Histoire [92]. Historical [71, 67, 128]. History [15, 69, 20, 95, 95, 92]. honorary [134, 106]. humanities [159]. hundred [48]. Hypothetical [221].

Idea [30, 164]. Ideal [188]. Ideal- [188]. Idealisation [30]. ideals [162]. Ideas [151]. Idées [151]. II [223, 223]. III [140]. IIIe [140]. im [91]. imperative [86]. Imperativsprache [86]. Importance [81]. Inaugural [38, 59, 192, 155]. Individual [178, 223]. Individuum [223, 178]. industrial [46]. industries [24]. Inleiding [1, 34]. Institute [220, 64, 119, 129, 130]. interests [132]. International [33, 57, 73, 121, 124, 136, 125, 156, 142, 167, 157, 154, 233, 155, 26, 13, 153, 43, 26]. Internationale [157, 154]. Internationaux [43]. interpretation [31]. intersubjectivity [196]. Introduction [36, 3, 4, 5, 6, 126, 55, 186, 194, 22, 34, 1, 34, 194]. Introductory [12, 14]. ISCS [13]. Isotype [185]. Ivanovich [94].

J [134, 107, 106]. Jaarwisseling [74]. Jac. [10]. Journal [43, 26]. journals [142, 167]. judging [84]. juridique [205].

Kalkuelengenetik [198]. knowledge [44, 28]. Korzybski [65].

L [134, 107]. L. [106]. Langage [187, 195, 196, 197]. langages [194]. Language [140, 197, 111, 199, 195, 202, 137, 66, 196, 63, 165, 187, 86]. Languages [151, 201, 194]. Langue [140]. Langues [151]. Laotse [189]. last [185]. Laws [85, 132, 62]. legal [205]. letter [15, 14]. level [63]. levels [196]. Life [141, 114]. Limes [30, 30]. Limes-Theory [30]. linguistic [200, 86, 122]. Linguistics’ [26, 157]. Linguistique [157, 26, 122]. linguistiques [200]. List [69]. livelihood [23]. Logic [215, 232, 147, 166, 145, 146, 72, 87, 148]. Logical [172, 111, 173, 186, 83, 163]. Logique [148, 232, 147]. Logische [163]. Lois [62]. lutte [117].

M [103]. M. [175]. main [69, 67, 93]. Mannoury [134, 106]. many [63]. many-level-structure [63]. Mass [91]. Massenpsychologie [91].

Mathematical [203]. mathematicians [102]. Mathematics [114, 108, 146, 187]. Mathématiques [187, 114]. meaning [212]. Meaningful [30]. mechanics [213]. Meeting [54, 50]. meilleure [25]. memoriam [228, 10]. mental [231]. metaphysical [209]. metaphysics [184]. Metarelation [198]. Method [90, 78, 207, 101, 160, 116, 205, 89]. Methode [160, 116, 78, 207, 101, 90, 205, 89]. Methoden [115]. methods [230, 210, 115]. microphysical [195]. microphysique [195]. microscopic [194, 195]. microscopiques [195, 194]. Mind [114, 132]. model [165]. Modell [165]. models [177]. Modern [219, 52, 53, 200, 152]. Moderne [52, 200]. moral [84]. Morals [162]. Morgan [144]. Moscow [15, 19]. mot [6]. Mots [151]. movement [60]. mutual [25]. mutuelle [25].

N. [228]. Nagel [31]. natural [177]. Nature [62, 24, 84]. Nederland [52]. néerlandaise [26]. Netherlands [125, 60, 39, 52, 53]. Neurath [8]. News [154]. niveaux [196]. non [210, 209]. non-metaphysical [209]. non-verbal [210]. notes [177]. Notice [58]. Nouvelles [154].

Object [116]. Objekt [116]. Objet [205]. Observables [187]. observations [123]. occasion [76, 106, 144]. oldest [179]. operation [13, 153]. operational [212]. origin [145]. Otto [8]. our [66, 28].

P. [218]. Paris [121]. part [37, 223]. partie [37]. Pays [39]. Pays-Bas [39]. peace [132]. pendant [41, 93]. phenomenalism [221]. philosopher [182]. philosophers [104, 118]. Philosophie [140, 232, 127, 214, 147, 90, 92, 233]. Philosophy [15, 119, 232, 44, 147, 90, 127, 92, 71, 214, 186, 213, 152, 110, 233, 140]. physical [186]. physics [200, 115]. Physik [115]. physique [200]. phytocenology [204]. pioneer [139]. plan [13]. Planned [188]. Planwirtschaft [188]. Plenary [50]. plus [179]. points [82]. Polish [170]. Politics [151]. Politique [151]. Positivism [176, 83, 175]. possible [210]. Postscript [174]. Pour [25, 26, 121]. pragmatic [216]. present [137]. President [15]. principal [95]. principaux [93]. principles [85]. Probability [30, 31, 229]. problem [82, 149, 49]. problème [82]. Problèmes [200, 93]. problems [95, 200, 93]. procedures [98]. Proceedings [140]. Prof. [107, 174]. Professor [134]. programme [192]. propos [77]. propositions [166]. proscription [180]. prospecten [99]. prospects [99]. Psychic [133]. psycho [200, 200]. Psycho-linguistic [200]. psycho-linguistiques [200]. Psychoanalysis [97]. Psychological [26, 157, 225, 111, 163]. psychologies [97]. Psychologique [157, 26]. psychologische [163]. psychology [91, 128, 103]. published [69]. Purpose [205].

quantum [213]. Quatrième [157]. que [214]. Quelques [82]. Qu’est [214]. Qu’est-ce [214]. qui [93]. quinquennat [93]. quinquennium [93]. racism [117]. racisme [117]. Rapport [127]. rays [183]. realism [176, 175]. Réalité [148]. Reality [148, 184, 188, 149, 45]. Realtypen [188]. reconstruction [91]. regard [108]. relations [161]. relativistic [108]. Remarks [83, 175, 189, 12]. renewal [76]. renouvellement [76]. reply [106]. Report [54, 213, 185, 127]. research [177, 19, 128]. researchers [20]. response [86, 66]. retrospect [144]. revelation [208]. Révész [86]. Révész’sche [86]. Review [11, 103, 135, 191, 32]. reviews [21, 143, 217, 226, 168]. revolt [218]. révolution [80, 80]. revue [26]. right [80]. role [177]. romain [179]. Roman [179]. Russian [69, 68, 104, 118].

’s [125, 193]. ’s-Graveland [125, 193]. S.J. [10]. Saturday [54]. Schlick [175]. Science [134, 121, 220, 141, 159, 90, 79, 127, 45, 197, 13, 16, 78, 101, 47, 184, 186, 205, 89, 117, 67, 98, 18, 150, 121, 129, 232, 130]. Sciences [127, 131, 177, 90, 89, 197, 93, 119, 232, 48, 93]. scientific [158]. scientists [104, 118]. Second [57, 54]. Section [40, 73, 15, 170, 56]. see [138]. semantical [208]. semantics [216, 64]. September [134, 106]. Sergei [94]. Series [124, 136]. seront [93]. service [13, 91]. Short [54]. signific [60]. Significal [57, 125, 156, 155]. significance [128]. Significant [122]. Significs [33, 125, 139, 203]. Signifique [122]. Sinnvolle [30]. six [27]. social [133, 177]. Societe [232]. Sociétés [140]. Societies [140]. Society [33, 232, 223, 178]. Sociobiology [109]. Sociological [171, 42, 180, 128]. Sociologie [43, 207, 41]. Sociologiques [171, 180, 42]. sociology [207, 41, 46, 43]. Some [177, 152, 123, 204, 82]. Soviet [50, 20, 19, 117, 49, 67, 18, 150, 102]. soviétique [117]. soziologische [128]. speech [61, 47, 163, 155]. Spirit [62]. Sprachbetrachtung [163]. Sprache [199, 165]. Sprachen [201]. Sprachgenetik [86]. Staat [178, 223]. stage [137]. Stalin [16]. standpoint [108]. State [223, 178]. statement [17]. statements [221, 210]. statics [219]. structure [66, 63]. students [73]. studied [93]. studies [71, 208, 171, 42]. Study [26, 33, 53, 20, 194, 204]. Summer [125, 157, 155, 57, 156]. survey [192]. symbolic [145, 72, 87]. symbols [219]. Synopsis [60]. Synthese [124, 136, 125, 121, 7, 26, 14]. Synthese-Series [124, 136]. Synthesis [121, 26]. system [66]. systematics [202]. Systèmes [195]. systems [195, 231].

Tao [189]. tasks [67]. teaching [159]. teachings [137]. technological [24]. Technology [23]. Teh [189]. Teil [223]. Ten [2, 112]. Tenth [50]. Teplov [103]. terms [204]. testament [179]. testing [210]. their [132]. theological [139]. Theorie [30]. Theory [30, 29, 229]. Third [125, 155]. Thirty [150]. thought [68, 47]. thoughts [199]. Tiefenpsychologie [128]. tot [34]. total [66]. totality [97]. totality-psychologies [97]. traditional [71]. Translation [15, 14]. travail [80]. turning [75]. Twenty [18, 48]. Twenty-eight [18]. Two [227, 48, 189]. Types [188].

U.S.S.R [119, 14]. U.S.S.R. [95, 48, 131]. Ueber [115, 222].

Understanding [222, 25]. Unesco [158]. Union [50]. Unit [160]. Unité [78, 101, 89]. Unity [129, 220, 130, 44, 89, 78, 101]. University [134]. upon [134].

Vavilov [94]. vegetation [204]. verbal [210]. verify [210]. Version [35, 134]. Verständlichkeit [201]. Verstehen [222]. Vice [15]. Vice-President [15]. vie [114]. view [82, 163]. Voies [154]. VOKS [15]. vue [82].

W [218]. Wahrscheinlichkeit [30]. war [180, 41]. ways [153]. Westendorp [228]. Wiederaufbaus [91]. Word [6, 161]. word-fact [161]. Words [151, 193]. work [158, 80, 185, 183]. works [69]. Writers [50].

XIe [233]. XIth [233].

year [76, 75, 17]. years [48, 27, 185, 18, 150].

References

Anonymous:1946:IDI

[1] Anonymous. Inleiding 1946. (Dutch) [Introduction 1946]. Synthese, 5 (1–2):3–5, May 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

     0964 (electronic). URL http://link.springer.com/article/10.1007/
     BF02273484;   http://link.springer.com/content/pdf/10.1007/BF02273484.
     pdf.

Anonymous:1946:TG

[2] Anonymous. Ten geleide 1940. Synthese, 5(1–2):6–8, May 1946. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

     http://link.springer.com/article/10.1007/BF02273485;   http:
     //link.springer.com/content/pdf/10.1007/BF02273485.pdf.

Anonymous:1946:Ia

[3] Anonymous. Introduction 1946. Synthese, 5(1–2):9–11, May 1946. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

     http://link.springer.com/article/10.1007/BF02273486;   http:
     //link.springer.com/content/pdf/10.1007/BF02273486.pdf.

Anonymous:1946:Ib

[4] Anonymous. Introduction 1940. Synthese, 5(1–2):12–14, May 1946. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

     http://link.springer.com/article/10.1007/BF02273487;   http:
     //link.springer.com/content/pdf/10.1007/BF02273487.pdf.

Anonymous:1946:Ic

[5] Anonymous. Introduction 1946. Synthese, 5(1–2):15–17, May 1946. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02273488;   http:
//link.springer.com/content/pdf/10.1007/BF02273488.pdf.

Anonymous:1946:MIF

[6] Anonymous. Un mot d’introduction (1940). (French) [A word of introduction (1940)]. Synthese, 5(1–2):18–20, May 1946. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF02273489;   http://link.
springer.com/content/pdf/10.1007/BF02273489.pdf.

Anonymous:1946:CSa

[7] Anonymous. The course of synthese. Synthese, 5(1–2):21–24, May 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02273490.

Anonymous:1946:ON

[8] Anonymous. Otto Neurath. Synthese, 5(1–2):24–25, May 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF02273491.

Boerma:1946:EED

[9] N. Westendorp Boerma. Empirie in de ethiek. (Dutch) [Empire in ethics]. Synthese, 5(1–2):26–35, May 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF02273492.

Mannoury:1946:MJV

[10] G. Mannoury. In memoriam Jac. Van Ginneken S.J. Synthese, 5(1–2): 35–37, May 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273493.

M:1946:BDB

[11] G. M. Boekbespreking. (Dutch) [Book review]. Synthese, 5(1–2):38– 43, May 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273494.

Anonymous:1946:IR

[12] Anonymous. Introductory remarks. Synthese, 5(1–2):44, May 1946. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02273495;   http:
//link.springer.com/content/pdf/10.1007/BF02273495.pdf.

Anonymous:1946:ISC

[13] Anonymous. The international science co-operation service plan (ISCS). Synthese, 5(1–2):46–51, May 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF02273496.

Valkov:1946:TIL

[14] Signed V. Valkov. Translation of an introductory letter for Synthese from His Excellency the Ambassador of the U.S.S.R. at the Hague. Synthese, 5(1–2):52, May 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273497;   http://link.springer.com/content/pdf/10.1007/BF02273497.
pdf.

Anonymous:1946:TLV

[15] Anonymous. Translation of a letter from the Vice-President of the Section History, Economics and Philosophy of the VOKS (Moscow). Synthese, 5 (1–2):52, May 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273498;   http://link.springer.com/content/pdf/10.1007/BF02273498.
pdf.

Anonymous:1946:SS

[16] Anonymous. Stalin on science. Synthese, 5(1–2):55–56, May 1946. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02273499.

Kapitsa:1946:NYS

[17] academician Peter Kapitsa. New year statement. Synthese, 5(1–2):56, May 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF02273500;
http://link.springer.com/content/pdf/10.1007/BF02273500.pdf.

Vavilov:1946:TEY

[18] Sergei Vavilov. Twenty-eight years of Soviet science. Synthese, 5(1–2): 57–59, May 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273501.

Lyashchenko:1946:MCS

[19] Academician P. Lyashchenko. Moscow — centre of Soviet research in economics. Synthese, 5(1–2):59–60, May 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF02273502.

Anonymous:1946:SRSa

[20] Anonymous. Soviet researchers study English history. Synthese, 5(1– 2):61, May 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273503;   http://link.springer.com/content/pdf/10.1007/BF02273503.
pdf.

Anonymous:1946:BR

[21] Anonymous. Book reviews. Synthese, 5(1–2):62, May 1946. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF02273504;   https://link.springer.
com/content/pdf/10.1007/BF02273504.pdf.

vanKleeck:1946:I

[22] Mary van Kleeck and Mary L. Fleddérus. Introduction. Synthese, 5 (1–2):63–64, May 1946. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273505;   https://link.springer.com/content/pdf/10.1007/
BF02273505.pdf.

Anonymous:1946:TL

[23] Anonymous. Technology and livelihood. Synthese, 5(1–2):64–65, May 1946. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF02273506.

Anonymous:1946:NTC

[24] Anonymous. The nature of technological changes in basic industries. Synthese, 5(1–2):65–70, May 1946. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF02273507.

Anonymous:1946:PMC

[25] Anonymous. Pour une meilleure comprehension mutuelle. (French) [For better mutual understanding]. Synthese, 5(1–2):71, May 1946.

CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF02273508;
https://link.springer.com/content/pdf/10.1007/BF02273508.pdf.

Anonymous:1946:CAR

[26] Anonymous. Le champ d’activité de la revue néerlandaise “Synthese” et du “Groupe international pour l’ètude de la linguistique psychologique”. (French) [The field of activity of the Dutch journal “Synthesis” and the ‘International Group for the Study of Psychological Linguistics’]. Synthese, 5(1–2):72–76, May 1946. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF02273509.

Neurath:1946:ASY

[27] Otto Neurath. After six years. Synthese, 5(1–2):77–82, May 1946. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF02273510.

Reach:1946:FOK

[28] K. Reach. The foundations of our knowledge. Synthese, 5(1–2):83–86, May 1946. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF02273511.

Geiringer:1946:FTE

[29] Hilda Geiringer. “Falsification” and theory of errors. Synthese, 5(1– 2):86–89, May 1946. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273512.

Strauss:1946:LTW

[30] M. Strauss. Ist Die Limes–Theorie Der Wahrscheinlichkeit Eine Sinnvolle Idealisation?. (German) [Is the Limes-theory of probability a meaningful idea?]. Synthese, 5(1–2):90–91, May 1946. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF02273513.

Beth:1946:DIP

[31] E. Beth. Discussion: On the interpretation of probability calculi: Ernest Nagel. Synthese, 5(1–2):92–95, May 1946. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF02273514.

Stebbing:1946:BR

[32] L. Susan Stebbing. Book review. Synthese, 5(1–2):95, May 1946. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF02273515;
https://link.springer.com/content/pdf/10.1007/BF02273515.pdf.

Anonymous:1946:EIS

[33] Anonymous. Editors: The International Society for the Study of Significs (Amsterdam). Synthese, 5(1–2):96, May 1946. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF02273516;   https://link.springer.com/
content/pdf/10.1007/BF02273516.pdf.

Anonymous:1946:ITI

[34] Anonymous. Inleiding tot een inleiding. (Dutch) [Introduction to an introduction]. Synthese, 5(1–2):97–100, May 1946. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF02273517.

Anonymous:1946:EV

[35] Anonymous. English version. Synthese, 5(1–2):101, May 1946. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF02273518;
https://link.springer.com/content/pdf/10.1007/BF02273518.pdf.

Anonymous:1946:II

[36] Anonymous. Introduction to an introduction. Synthese, 5(1–2):102– 104, May 1946. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273519;   https://link.springer.com/content/pdf/10.1007/
BF02273519.pdf.

Anonymous:1946:AGG

[37] Anonymous. Algemeen gedeelte — general part — partie générale. Synthese, 5(3–4):107, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF02273734.

Anonymous:1946:IAa

[38] Anonymous. Inaugural address. Synthese, 5(3–4):108–115, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02273735.

Huberthrinin:1946:AFP

[39] Huberthrinin. Ambassade de France aux Pays-Bas. (French) [Embassy of France in The Netherlands]. Synthese, 5(3–4):116, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF02273736.

Anonymous:1946:SF

[40] Anonymous. Section française. (French) [French section]. Synthese, 5 (3–4):117, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273737.

Polin:1946:SFP

[41] Raymond Polin. La sociologie française pendant la guerre. (French) [French sociology during the war]. Synthese, 5(3–4):117–129, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02273738.

Levy-Bruhl:1946:CES

[42] Henri Lévy-Bruhl. Le centre d’études sociologiques. (French) [The Center for Sociological Studies]. Synthese, 5(3–4):130–132, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF02273739.

Anonymous:1946:CIS

[43] Anonymous. Cahiers internationaux de sociologie. (French) [International Journal of Sociology]. Synthese, 5(3–4):133, July 1946. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF02273740.

Emmet:1946:PUK

[44] Dorothy M. Emmet. Philosophy and “the unity of knowledge”. Synthese, 5(3–4):134–137, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF02273741.

Polanyi:1946:SR

[45] Michael Polanyi. Science and reality. Synthese, 5(3–4):137–150, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02273742.

vanKleeck:1946:TIS

[46] Mary van Kleeck. Towards an industrial sociology. Synthese, 5(3–4): 151–155, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273743.

Dobrogayev:1946:SST

[47] S. Dobrogayev. The science of speech and thought. Synthese, 5(3–4): 156–161, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273744.

Komarov:1946:THT

[48] V. L. Komarov. Two hundred and twenty years of the U.S.S.R. Academy of Sciences. Synthese, 5(3–4):161–166, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF02273745.

Traehtenberg:1946:SEP

[49] Academician I. A. Traehtenberg. I. Soviet economists on the problem of cycles and crises. Synthese, 5(3–4):167–169, July 1946. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF02273746.

Anonymous:1946:TPM

[50] Anonymous. The Tenth Plenary Meeting of the Union of Soviet Writers. Synthese, 5(3–4):170–173, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF02273747.

Anonymous:1946:NB

[51] Anonymous. New books. Synthese, 5(3–4):173, July 1946. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF02273748.

Kruseman:1946:AMB

[52] W. M. Kruseman. Aspecten der moderne biologie in Nederland. (German) [Aspects of modern biology in The Netherlands]. Synthese, 5(3–4): 174–181, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273749.

Kruseman:1946:ASM

[53] W. M. Kruseman. Aspects of the study of modern biology in The Netherlands. Synthese, 5(3–4):181–182, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF02273750.

Anonymous:1946:SRSb

[54] Anonymous. Short report of the second meeting held on Saturday Afternoon February 1st 1941. Synthese, 5(3–4):183–186, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF02273751.

Beth:1946:I

[55] E. W. Beth. Introduction. Synthese, 5(3–4):187, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF02273752.

Anonymous:1946:BS

[56] Anonymous. British section. Synthese, 5(3–4):188, July 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF02273753.

Anonymous:1946:SISa

[57] Anonymous. Second international significal Summer conference 1946. Synthese, 5(5–6):191, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF02274377.

Anonymous:1946:N

[58] Anonymous. Notice. Synthese, 5(5–6):192, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF02274378.

Anonymous:1946:IAb

[59] Anonymous. Inaugural address. Synthese, 5(5–6):193–200, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF02274379.

Brouwer:1946:SSM

[60] L. E. J. Brouwer. Synopsis of the signific movement in The Netherlands. Synthese, 5(5–6):201–208, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF02274380.

Anonymous:1946:CSb

[61] Anonymous. Concluding speech. Synthese, 5(5–6):209–212, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF02274381.

Mercier:1946:LNC

[62] André Mercier. Lois de la nature et constructions de l’esprit. (French) [Laws of nature and constructions of the spirit]. Synthese, 5(5–6):213– 220, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02274382.

Waismann:1946:MLS

[63] Friedrich Waismann. The many-level-structure of language. Synthese, 5(5– 6):221–229, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF02274383.

Anonymous:1946:IGS

[64] Anonymous. The Institute of General Semantics. Synthese, 5(5–6):230– 231, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02274384.

Reesinck:1946:K

[65] J. J. M. Reesinck. On Korzybski. Synthese, 5(5–6):231–236, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF02274385.

Pritchard:1946:ELS

[66] Earle A. Pritchard. Everyday language and the structure of our total response system. Synthese, 5(5–6):236–242, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF02274386.

Udaltsov:1946:MTS

[67] A. Udaltsov. The main tasks of Soviet historical science. Synthese, 5(5–6): 243–244, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF02274387.

Anonymous:1946:DRE

[68] Anonymous. The development of Russian economic thought. Synthese, 5(5–6):245–251, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF02274388.

Anonymous:1946:BLM

[69] Anonymous. Bibliography. List of main works on Russian history published between 1940 and 1945. Synthese, 5(5–6):251–253, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02274389.

Anonymous:1946:HB

[70] Anonymous. Henri Bergson. Synthese, 5(5–6):254–257, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02274390.

Beth:1946:HST

[71] Evert W. Beth. Historical studies in traditional philosophy. Synthese, 5(5– 6):258–270, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF02274391.

Meyer:1946:DSLa

[72] H. Meyer. On definitions in symbolic logic. Synthese, 5(5–6):271–276, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02274392.

Anonymous:1946:SISb

[73] Anonymous. Section: International students contact. Synthese, 5(5–6): 277–280, September 1946. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF02274393.

Anonymous:1946:BJD

[74] Anonymous. Bij de jaarwisseling. (Dutch) [At the annual exchange]. Synthese, 5(7–8):283–284, November 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF02273754.

Anonymous:1946:TY

[75] Anonymous. At the turning of the year. Synthese, 5(7–8):284–285, November 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273755.

Anonymous:1946:ODR

[76] Anonymous. À l’occasion du renouvellement de l’année. (French) [On the occasion of the renewal of the year]. Synthese, 5(7–8):285–286, November 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF02273756.

Anonymous:1946:AP

[77] Anonymous. Avant-propos. (French) [Foreword]. Synthese, 5(7–8):287, November 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273757.

Anonymous:1946:UMS

[78] Anonymous. Unité et méthode dans la science. (French) [Unity and method in science]. Synthese, 5(7–8):288–295, November 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF02273758.

Morris:1946:SD

[79] Charles W. Morris. Science and discourse. Synthese, 5(7–8):296–308, November 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273759.

Bourgin:1946:DAT

[80] Georges Bourgin. Le “droit au travail” et la révolution de 1848 en France. (French) [The “right to work” and the Revolution of 1848 in France]. Synthese, 5(7–8):309–315, November 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF02273760.

Emmet:1946:I

[81] Dorothy M. Emmet. Importance. Synthese, 5(7–8):316–320, November 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF02273761.

Bernays:1946:QPV

[82] P. Bernays. Quelques points de vue concernant le problème de l’évidence. (French) [Some points of view concerning the problem of the evidence]. Synthese, 5(7–8):321–326, November 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF02273762.

Kazemier:1946:RLP

[83] Brugt H. Kazemier. Remarks on logical positivism. Synthese, 5(7–8): 327–332, November 1946. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF02273763.

Boerma:1946:NMJ

[84] N. Westendorp Boerma. On the nature of moral judging. Synthese, 5(7– 8):332–338, November 1946. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF02273764.

Clay:1946:LP

[85] J. Clay. Laws and principles. Synthese, 5(7–8):338–348, November 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02273765.

Mannoury:1946:RSA

[86] G. Mannoury. Die Révész’sche Sprachgenetik Aufforderungsfunktion und Imperativsprache. (German) [The Révész linguistic genetics response function and imperative language]. Synthese, 5(7–8):348–351, November 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF02273766.

Meyer:1946:DSLb

[87] H. Meyer. On definitions in symbolic logic. Synthese, 5(7–8):353–361, November 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02273767.

Anonymous:1946:E

[88] Anonymous. Erratum. Synthese, 5(7–8):361, November 1946. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF02273768;   http://link.
springer.com/content/pdf/10.1007/BF02273768.pdf.

Reymond:1947:UMS

[89] Arnold Reymond. De l’unité et de la méthode dans les sciences. (French) [Of unity and method in science]. Synthese, 5(9–10):365–378, January 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF02267708.

Gonseth:1947:BMP

[90] Ferdinand Gonseth. Sur les buts et la méthode de la philosophie des sciences. (French) [On the goals and method of the philosophy of science]. Synthese, 5(9–10):379–389, January 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF02267709.

Reiwald:1947:MID

[91] Paul Reiwald. Die Massenpsychologie im Dienste des Wiederaufbaus. (German) [Mass psychology in the service of reconstruction]. Synthese, 5(9–10):390–394, January 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF02267710.

Vychinski:1947:HPF

[92] P. Vychinski. Histoire de la philosophie. (French) [History of philosophy]. Synthese, 5(9–10):395–410, January 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF02267711.

Vavilov:1947:PPQ

[93] l’académicien S. I. Vavilov. Les principaux problèmes qui seront étudiés par l’académie des sciences pendant le quinquennat en cours. (French) [The main problems to be studied by the Academy of Sciences during the current quinquennium]. Synthese, 5(9–10):411–421, January 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02267712.

Gorelov:1947:SIV

[94] S. N. Gorelov. Sergei Ivanovich Vavilov. Synthese, 5(9–10):422–424, January 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02267713.

Anonymous:1947:HPP

[95] Anonymous. History I. The principal problems of history in the U.S.S.R. Synthese, 5(9–10):425–429, January 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF02267714.

Anonymous:1947:NBa

[96] Anonymous. New book. Synthese, 5(9–10):429, January 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF02267715.

Holstijn:1947:PTP

[97] A. J. Westerman Holstijn. Psychoanalysis and totality-psychologies. Synthese, 5(9–10):431–440, January 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF02267716.

VanDantzig:1947:GPE

[98] D. Van Dantzig. General procedures of empirical science. Synthese, 5(9– 10):441–455, January 1947. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF02267717.

Anonymous:1947:APD

[99] Anonymous. Aspecten en prospecten. (Dutch) [Aspects and prospects]. Synthese, 5(11–12):461–471, March 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF02268413.

Anonymous:1947:CBF

[100] Anonymous. Congrès de Berne. (French) [Congress of Berne]. Synthese, 5 (11–12):473–485, March 1947. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF02268414.

Clay:1947:UMS

[101] J. Clay. Unité et méthode dans la science. (French) [Unity and method in science]. Synthese, 5(11–12):486–498, March 1947. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF02268415.

Vinogradov:1947:SM

[102] Academician I. Vinogradov. Soviet mathematicians. Synthese, 5(11–12): 499–503, March 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02268416.

Smirnov:1947:BRB

[103] A. Smirnov. Book review: B. M. Teplov. “Psychology”. Synthese, 5(11– 12):503–505, March 1947. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF02268417.

Frolov:1947:RSP

[104] Prof. Y. P. Frolov. Russian scientists and philosophers. Synthese, 5(11– 12):506–510, March 1947. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF02268418.

Mannoury:1947:E

[105] G. Mannoury. Editorial. Synthese, 5(11–12):511–513, March 1947. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02268419.

Anonymous:1947:GMR

[106] Anonymous. G. Mannoury’s reply to the address delivered by L. E. J. Brouwer on the occasion of his honorary doctorate on September 16th, 1946. Synthese, 5(11–12):514–515, March 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF02268420.

Anonymous:1947:CGD

[107] Anonymous. Commemoration of the graduation day of Prof. Dr. L. E. J. Brouwer, February 19, 1907. Synthese, 5(11–12):516–518, March 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02268421.

Anonymous:1947:RSR

[108] Anonymous. The relativistic standpoint with regard to the foundation of mathematics. Synthese, 5(11–12):519–521, March 1947. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF02268422.

Anonymous:1947:S

[109] Anonymous. Sociobiology. Synthese, 5(11–12):522–525, March 1947. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02268423.

tenHave:1947:EHP

[110] T. T. ten Have. Essentials of Heymans’ philosophy. Synthese, 5(11–12): 526–541, March 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02268424.

Beth:1947:LPA

[111] Evert W. Beth. Logical and psychological aspects in the consideration of language. Synthese, 5(11–12):542–544, March 1947. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF02268425.

Anonymous:1947:TG

[112] Anonymous. Ten geleide. (Dutch) [Guided]. Synthese, 6(1–2):5–10, May 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00414088.

Anonymous:1947:CBA

[113] Anonymous. Congrès de Berne (11, 12 et 13 avril 1946). (French) [Congress of Berne (11, 12 and 13 April 1946)]. Synthese, 6(1–2):11, May 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00414089.

Wavre:1947:VEM

[114] Rolin Wavre. La vie de l’esprit dans les mathématiques. (French) [The life of the mind in mathematics]. Synthese, 6(1–2):12–24, May 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00414090.

Konig:1947:MPG

[115] Hans König. Ueber die Methoden der Physik. (German) [On the methods of physics]. Synthese, 6(1–2):25–43, May 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00414091.

Lehmann:1947:OMB

[116] F. E. Lehmann. Objekt und Methode in der Biologie. (German) [Object and method in biology]. Synthese, 6(1–2):44–56, May 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00414092.

Roguinski:1947:SSL

[117] J. Roguinski, Licencié ès-sciences biologiques. La science soviétique dans la lutte contre le racisme. (French) [Soviet science in the fight against racism]. Synthese, 6(1–2):59–68, May 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00414093.

Maximov:1947:RSP

[118] Prof. A. A. Maximov. Russian scientists and philosophers. Synthese, 6(1–2):69–71, May 1947. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00414094.

Anonymous:1947:IPU

[119] Anonymous. In the Institute of Philosophy, U.S.S.R. Academy of Sciences. Synthese, 6(1–2):71, May 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00414095.

Anonymous:1947:NBb

[120] Anonymous. New books. Synthese, 6(1–2):72–73, May 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00414096.

Anonymous:1947:CFP

[121] Anonymous. Communication de la fondation “Pour La Science” Centre International de Synthèse (Paris). (French) [Communication from the Foundation “For Science” International Center for Synthesis (Paris)]. Synthese, 6(1–2):77, May 1947. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00414097.

Stutterheim:1947:SLF

[122] C. F. P. Stutterheim. Signifique et linguistique. (French) [Significant and linguistic]. Synthese, 6(1–2):81–88, May 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00414098.

Monnich:1947:SOC

[123] C. W. Mönnich. Some observations on the character of the Christian dogma. Synthese, 6(1–2):89–98, May 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00414099.

Anonymous:1947:SSIa

[124] Anonymous. The synthese-series and the international book-committee. Synthese, 6(1–2):101–106, May 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00414100.

Anonymous:1947:TIS

[125] Anonymous. The Third International Significal Summer Conference (’s-Graveland, Netherlands, August 1947), Significs and “Synthese”. Synthese, 6(3–4):113–121, July 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00490176.

Anonymous:1947:I

[126] Anonymous. Introduction. Synthese, 6(3–4):122–127, July 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00490177.

Piaget:1947:DRS

[127] Jean Piaget. Du rapport des sciences avec la philosophie. (French) [From the report of science to philosophy]. Synthese, 6(3–4):130–150, July 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00490178.

Walter:1947:BTG

[128] Emil J. Walter. Die Bedeutung der Tiefenpsychologie für die geschichtliche und soziologische Forschung. (German) [The significance of depth psychology for historical and sociological research]. Synthese, 6(3–4):152– 156, July 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00490179.

Anonymous:1947:IUS

[129] Anonymous. Institute for the Unity of Science. Synthese, 6(3–4):158– 159, July 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00490180.

Frank:1947:IUS

[130] Philipp Frank. The Institute for the Unity of Science. Synthese, 6(3–4): 160–167, July 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00490181.

Vavilov:1947:EAS

[131] Sergei Vavilov. Elections to the Academy of Sciences of the U.S.S.R. Synthese, 6(3–4):170–173, July 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00490182.

West:1947:FLM

[132] Ranyard West. Fixed laws of the mind and their control in the interests of peace. Synthese, 6(3–4):176–181, July 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00490183.

Claremont:1947:PCS

[133] Claude A. Claremont. Psychic conditions of social happiness. Synthese, 6(3–4):182–188, July 1947. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00490184.

Anonymous:1947:ADS

[134] Anonymous. Address delivered on September 16th, 1946, at the University of Amsterdam by Professor L. E. J. Brouwer on the conferment upon Professor G. Mannoury of the honorary degree of Doctor of Science (English version). Synthese, 6(3–4):190–194, July 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00490185.

Beth:1947:BR

[135] Evert W. Beth. Book review. Synthese, 6(3–4):195, July 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00490186.

Anonymous:1947:SSIb

[136] Anonymous. The synthese-series and the international book-committee. Synthese, 6(3–4):198–199, July 1947. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00490187.

Meshchaninov:1948:NTC

[137] Academician I. Meshchaninov. New teachings concerning language at its present stage of development. Synthese, 6(5–6):204–212, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00590982.

denHollander:1948:SU

[138] A. N. J. den Hollander. As others see us. Synthese, 6(5–6):214–237, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00590983.

Mannoury:1948:PTS

[139] G. Mannoury. A pioneer of theological significs. Synthese, 6(5–6):239– 241, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00590984.

Beth:1948:ADI

[140] E. W. Beth. Actes du IIIe Congrès des Sociétés de Philosophie de Langue française. (French) [Proceedings of the III Congress of the Societies of French Language Philosophy]. Synthese, 6(5–6):244–245, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00590985.

Choynowski:1948:LS

[141] Mieczyslaw Choynowski. Life of science. Synthese, 6(5–6):248–251, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00590986.

Anonymous:1948:IJa

[142] Anonymous. International journals. Synthese, 6(5–6):254–256, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00590987.

B:1948:BR

[143] W. B. and V. Book reviews. Synthese, 6(5–6):257–262, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00590988.

Anonymous:1948:ROB

[144] Anonymous. A retrospect on the occasion of the Boole–De Morgan centenary. Synthese, 6(7–8):267, July 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00590724.

Beth:1948:OGS

[145] E. W. Beth. The origin and growth of symbolic logic. Synthese, 6(7–8): 268–274, July 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00590725.

Heyting:1948:FLM

[146] Arend Heyting. Formal logic and mathematics. Synthese, 6(7–8):275– 282, July 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00590726.

Feys:1948:LFP

[147] Robert Feys. Logique formalisée et philosophie. (French) [Formal logic and philosophy]. Synthese, 6(7–8):283–298, July 1948. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00590727.

Destouches:1948:LRF

[148] Jean-Louis Destouches. Logique et réalité. (French) [Logic and reality]. Synthese, 6(7–8):300–304, July 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00590728.

Clay:1948:PR

[149] J. Clay. The problem of reality. Synthese, 6(7–8):305–316, July 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00590729.

Vavilov:1948:TYS

[150] Sergei Vavilov. Thirty years of Soviet science. Synthese, 6(7–8):318– 329, July 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00590730.

Bourgin:1948:IML

[151] Georges Bourgin. Idées, mots, langues en politique et en diplomatie. (French) [Ideas, words, languages in politics and diplomacy]. Synthese, 6(7–8):332–335, July 1948. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00590731.

Meyer:1948:SAM

[152] Herman Meyer. Some aspects of modern philosophy. Synthese, 6(7–8): 338–355, July 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00590732.

Anonymous:1948:NWI

[153] Anonymous. New ways in international co-operation. Synthese, 6(9–12): 361–365, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00590705.

Anonymous:1948:VNC

[154] Anonymous. Voies nouvelles dans la coopération internationale. (French) [News channels in international cooperation]. Synthese, 6(9–12):366–370, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00590706.

deRanitz:1948:IST

[155] Jhr. Dr. C. J. A. de Ranitz. Inaugural speech (Third International Significal Summer Conference). Synthese, 6(9–12):371, September 1948. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00590707;   http:
//link.springer.com/content/pdf/10.1007/BF00590707.pdf.

Anonymous:1948:FIS

[156] Anonymous. Fourth international significal Summer conference. Synthese, 6(9–12):372–374, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00590708.

Anonymous:1948:QCE

[157] Anonymous. La quatrième conférence d’été internationale de linguistique psychologique. (French) [The Fourth International Summer Conference on Psychological Linguistics]. Synthese, 6(9–12):375–377, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00590709.

Anonymous:1948:USW

[158] Anonymous. On Unesco’s scientific work. Synthese, 6(9–12):378–380, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00590710.

Frank:1948:STH

[159] Philipp Frank. Science teaching and the humanities. Synthese, 6(9–12): 382–410, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00590711.

Fischer:1948:EMB

[160] Hans Fischer. Einheit und Methode in der Biologie. (German) [Unit and method in biology]. Synthese, 6(9–12):412–418, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00590712.

Esser:1948:WFR

[161] P. H. Esser. On word-fact relations. Synthese, 6(9–12):420–425, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00590713.

D:1948:MEI

[162] M. Lietaert Peerbolte M. D. Morals, ethics, ideals. Synthese, 6(9–12): 426–457, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00590714.

Waismann:1948:LPA

[163] Friedrich Waismann. Logische und psychologische Aspekte in der Sprachbetrachtung. (German) [Logical and psychological aspects in the speech view]. Synthese, 6(9–12):460–475, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00590715.

Daan:1948:IF

[164] Albert Daan. The idea of freedom. Synthese, 6(9–12):476–486, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00590716.

Denk:1948:SME

[165] F. Denk. Sprache, Modell und Exaktheit. (German) [Language, model and accuracy]. Synthese, 6(9–12):487–494, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:

//link.springer.com/article/10.1007/BF00590717.

Ridder:1948:LP

[166] J. Ridder. Logic of propositions. Synthese, 6(9–12):496–502, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00590718.

Anonymous:1948:IJb

[167] Anonymous. International journals. Synthese, 6(9–12):504–507, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic).  URL http://link.springer.com/article/10.1007/
BF00590719.

Meyer:1948:BR

[168] H. Meyer and E. W. Beth. Book reviews. Synthese, 6(9–12):510–514, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00590720.

Anonymous:1948:NB

[169] Anonymous. New books. Synthese, 6(9–12):515, September 1948. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00590721;   http:
//link.springer.com/content/pdf/10.1007/BF00590721.pdf.

Anonymous:1948:PS

[170] Anonymous. Polish section. Synthese, 6(9–12):518, September 1948. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00590722;   http:
//link.springer.com/content/pdf/10.1007/BF00590722.pdf.

Anonymous:1948:CES

[171] Anonymous. Centre d’Études sociologiques. (French) [Center for Sociological Studies]. Synthese, 6(9–12):519–520, September 1948. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00590723.

Anonymous:1948:LE

[172] Anonymous. Logical empiricism. Synthese, 7(1):457, January 1948-1949. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00540039.

Frank:1948:LE

[173] Phillipp Frank. Logical empiricism I. Synthese, 7(1):458–465, January 1948-1949. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00540040.

Clay:1948:PPC

[174] J. Clay. Postscript by Prof. clay. Synthese, 7(1):465, January 1948-1949. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00540041.

Rynin:1948:RMS

[175] D. Rynin. Remarks on M. Schlick’s essay “Positivism and realism”. Synthese, 7(1):466–477, January 1948-1949. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00540042.

Schlick:1948:PR

[176] M. Schlick. Positivism and realism. Synthese, 7(1):478–505, January 1948-1949. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00540043.

Deutsch:1948:SNR

[177] Karl Wolfgang Deutsch. Some notes on research on the role of models in the natural and social sciences. Synthese, 7(1):506–533, January 1948-1949. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00540044.

Walter:1948:IGS

[178] Emil J. Walter. Individuum, Gesellschaft und Staat. (German) [Individual, society and state]. Synthese, 7(1):536–549, January 1948-1949. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00540045.

Levy-Bruhl:1948:PAT

[179] Henri Lévy-Bruhl. Le plus ancien testament romain. (French) [The oldest Roman testament]. Synthese, 7(1):552–553, January 1948-1949. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00540046.

Kopelmanas:1948:CSP

[180] L. Kopelmanas. Les conditions sociologiques de la proscription de la guerre. (French) [The sociological conditions of the proscription of war]. Synthese, 7(1):554–557, January 1948-1949. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00540047.

Clay:1948:CE

[181] J. Clay. The concept of evolution. Synthese, 7(1):560–574, January 1948-1949. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00540048.

Mannoury:1949:CP

[182] G. Mannoury. Clay as a philosopher. Synthese, 8(1):1–2, January 1949– 1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485882.

denHartog:1949:CWC

[183] H. den Hartog. Clay’s work on cosmic rays. Synthese, 8(1):3–5, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485883.

Ducasse:1949:RSM

[184] C. J. Ducasse. Reality, science, and metaphysics. Synthese, 8(1):9–21, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485884.

Neurath:1949:RLY

[185] Marie L. Neurath. Report on the last years of isotype work. Synthese, 8(1): 22–27, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485885.

Frank:1949:IPP

[186] Philipp Frank. Introduction to the philosophy of physical science, on the basis of logical empiricism. Synthese, 8(1):28–45, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485886.

Servien:1949:LMO

[187] Pius Servien. Langage, mathématiques et observables. (French) [Language, mathematics and observables]. Synthese, 8(1):49–53, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485887.

Walter:1949:IRP

[188] Emil J. Walter. Ideal- Und Realtypen Der Planwirtschaft. (German) [Ideal and reality types of planned economy]. Synthese, 8(1):54–64, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485888.

Stern:1949:RTC

[189] Axel L. Stern. Remarks on two chapters of Laotse’s Tao Teh Ching. Synthese, 8(1):65–69, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485889.

Meyer:1949:C

[190] Justus Meyer. Contingency. Synthese, 8(1):73–90, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485890.

Martin:1949:BRa

[191] Norman M. Martin. Book review. Synthese, 8(1):93–94, January 1949– 1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485891.

Esser:1949:IAG

[192] P. H. Esser. Inaugural address and general survey of the programme. Synthese, 8(1):101–107, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485892.

Schimmelpenninck:1949:WGB

[193] Count Schimmelpenninck. Words of greeting by the Burgomaster of ’s-Graveland. Synthese, 8(1):108, January 1949–1951. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485893.

Mannoury:1949:IEL

[194] G. Mannoury. Introduction à l’étude de langages microscopiques. (French) [Introduction to the study of microscopic languages]. Synthese, 8(1):111– 112, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485894.

Destouches-Fevrier:1949:SML

[195] Paulette Destouches-Fevrier. Systèmes microscopiques et langage microphysique. (French) [Microscopic systems and microphysical language]. Synthese, 8(1):113–119, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485895.

Rougier:1949:DNI

[196] Louis Rougier. Différents niveaux d’intersubjectivité du langage. (French) [Different levels of intersubjectivity of language]. Synthese, 8(1):120–129, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485896.

Servien:1949:LSF

[197] Pius Servien. Le langage des sciences. (French) [The language of science]. Synthese, 8(1):130–133, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485897.

Szekely:1949:MKG

[198] David L. Szekely. Metarelation und Kalkuelengenetik. (German) [Metarelation and calculus genetics]. Synthese, 8(1):134–147, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485898.

Denk:1949:BCS

[199] Franz Denk. Das A–B–C der Sprache und das A–B–C der Gedanken. (German) [The A–B–C of language and the A–B–C of thoughts]. Synthese, 8(1):148–154, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485899.

Destouches:1949:PPL

[200] Jean-Louis Destouches. Problèmes psycho-linguistiques en physique moderne. (French) [Psycho-linguistic problems in modern physics]. Synthese, 8(1):155–166, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485900.

Britzelmayr:1949:VFS

[201] Wilhelm Britzelmayr. Über die Verständlichkeit formalisierter Sprachen. (German) [On the comprehensibility of formalized languages]. Synthese, 8(1):167–171, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485901.

Kruseman:1949:GLB

[202] W. M. Kruseman. Gradation of language in biological systematics. Synthese, 8(1):175–181, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485902.

Rapoport:1949:MBC

[203] Anatol Rapoport. Mathematical biophysics, cybernetics and significs. Synthese, 8(1):182–193, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485903.

Westhoff:1949:ASC

[204] V. Westhoff. An analysis of some concepts and terms in vegetation study or phytocenology. Synthese, 8(1):194–206, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485904.

Levy-Bruhl:1949:OMS

[205] Henri Levy-Bruhl. Objet et méthode de la science juridique. (French) [Purpose and method of legal science]. Synthese, 8(1):209–212, January 1949– 1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485905.

Boerma:1949:CG

[206] N. Westendorp Boerma. Collective guilt. Synthese, 8(1):213–221, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485906.

Bourgin:1949:MES

[207] Georges Bourgin. La méthode empirique en sociologie. (French) [The empirical method in sociology]. Synthese, 8(1):222–224, January 1949– 1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485907.

Tonnessen:1949:FAR

[208] Herman Tönnessen. The fight against revelation in semantical studies. Synthese, 8(1):225–234, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485908.

Stern:1949:DNM

[209] Axel L. Stern. In defence of non-metaphysical ethics. Synthese, 8(1): 235–237, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485909.

D:1949:IPV

[210] P. H. Esser M. D. Is it possible to verify statements by non-verbal testing methods? Synthese, 8(1):238–242, January 1949–1951. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485910.

Esser:1949:FA

[211] P. H. Esser. Final address. Synthese, 8(1):243–246, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485911.

Bridgman:1949:OAM

[212] P. W. Bridgman. The operational aspect of meaning. Synthese, 8(1): 251–259, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485912.

Margenau:1949:RRD

[213] Henry Margenau and John Compton. Report on recent developments in the philosophy of quantum mechanics. Synthese, 8(1):260–271, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485913.

Ducasse:1949:QCQ

[214] C. J. Ducasse. Qu’est-ce que la philosophie?. (French) [What is philosophy?]. Synthese, 8(1):272–286, January 1949–1951. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485914.

Ajdukiewicz:1949:LE

[215] Kazimierz Ajdukiewicz. Logic and experience. Synthese, 8(1):289–299, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485915.

Grzegorczyk:1949:PFS

[216] Andrzej Grzegorczyk. The pragmatic foundations of semantics. Synthese, 8(1):300–324, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485916.

Kotarbinski:1949:BR

[217] Tadeusz Kotarbinski. Book reviews. Synthese, 8(1):325–327, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485917.

Bernstein:1949:PWB

[218] Jeremy Bernstein. P. W. Bridgman, in revolt against formalism. Synthese, 8(1):331–341, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485918.

Frank:1949:MAS

[219] Josef Frank. Modern architecture and the symbols of statics. Synthese, 8(1):342–349, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485919.

Anonymous:1949:IUS

[220] Anonymous. The Institute for the Unity of Science Essay Contest. Synthese, 8(1):350–351, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485920.

Brown:1949:HSP

[221] Robert Brown and John Watling. Hypothetical statements in phenomenalism. Synthese, 8(1):355–366, January 1949–1951. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485921.

Schachter:1949:VG

[222] Josef Schächter. Ueber das Verstehen. (German) [On understanding]. Synthese, 8(1):367–384, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485922.

Walter:1949:ITI

[223] Emil J. Walter. (II. Teil) Individuum, Gesellschaft und Staat. (German) [(Part II) Individual, society and state]. Synthese, 8(1):385–399, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485923.

Levy-Bruhl:1949:A

[224] Henri Levy-Bruhl. Assemblée. (French) [Assembly]. Synthese, 8(1):400– 406, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485924.

Tatarkiewicz:1949:PH

[225] Władysław Tatarkiewicz. Psychological hedonism. Synthese, 8(1):409– 425, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485925.

Martin:1949:BRb

[226] Norman M. Martin and B. H. Kazemier. Book reviews. Synthese, 8(1): 429–431, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485926.

Bergmann:1949:TCE

[227] Gustav Bergmann. Two cornerstones of empiricism. Synthese, 8(1):435– 452, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485927.

Anonymous:1949:MWB

[228] Anonymous. In memoriam N. Westendorp Boerma. Synthese, 8(1):453– 455, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485928.

vanDantzig:1949:CFP

[229] D. van Dantzig. Carnap’s foundation of probability theory. Synthese, 8(1): 459–470, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485929.

Clay:1949:CM

[230] J. Clay. Combination of methods. Synthese, 8(1):471–479, January 1949– 1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485930.

Kazemier:1949:FSM

[231] B. H. Kazemier. Formal systems and mental activity. Synthese, 8(1): 480–489, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485931.

Destouches:1949:SFL

[232] J. L. Destouches. Societe Française de Logique et de Philosophie des Sciences. (French) [French Society of Logic and Philosophy of Science]. Synthese, 8(1):490, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485932.

Anonymous:1949:XCI

[233] Anonymous. Le XIe Congres International de Philosophie. (French) [The XIth International Congress of Philosophy]. Synthese, 8(1):491, January 1949–1951. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485933.